OpenStudy (yessenia27):

(02.03 HC) Grant has fraction 2 over 3 of a cup of ice cream in a bowl. He then removes two fraction 1 over 6 cups of ice cream to share with his little brother. Which of the following best explains the amount of ice cream left in Grant's bowl? fraction 26, because fraction 2 over 3 minus 2 times fraction 1 over 6 equals fraction 2 over 3 minus fraction 3 over 6 fraction 2 over 6, because fraction 2 over 3 minus 2 times fraction 1 over 6 equals fraction 2 over 3 minus fraction 2 over 6 fraction 1 over 6, because fraction 2 over 3 minus 2 times fraction 1 over 6 equals fraction 2 ove

OpenStudy (yanasidlinskiy):

2/3 - 2/6 = 4/6 - 2/6 = 2/6 which rounds to 1/3....so basically, when rounded, he has 1/3 cup of ice cream left